The domain of definition of the function y (x) is given by the equation 2x+ 2y = 2 is :
(A) 0< x≤1                                                  (B) 0≤x≤1
(C) -∞ < x ≤0                                              (D) -∞< x<1

Grade:12

1 Answers

AskIITian Expert Priyasheel - IITD
14 years ago

2^x +2^y =2, or, 2^y = 2 - 2^x

Take log with base 2 on both sides,

y=log(2-2^x),

For y to be real, 2 - 2^x > 0,

i.e., 2 > 2^x,

or, x < 1.

ANS is (D).